# All-Sky Kinematics of the Distant Halo: The Reflex Response to the LMC

### v1

@@ 1 Introduction
The LMC and SMC (hereafter the ‘Clouds’) are by far the Milky Way’s (MW’s) most prominent satellite galaxies, and were likely the first external galaxies discovered (e.g., Dennefeld, 2020). There is growing evidence that the Clouds form a binary pair — with the LMC being 10$\approx$ 10$\times$ 10 × more massive than the SMC — and are on their first orbital infall towards the Milky Way (Kallivayalil et al., 2006; Besla et al., 2007, 2010; Kallivayalil et al., 2013, but see Vasiliev 2024). Multiple independent measurements estimate the total mass of the LMC to be between $1.0-2.5\times 10^{11}\,M_{\odot}$ (Peñarrubia et al., 2016; Erkal et al., 2019b, 2020; Vasiliev et al., 2021; Shipp et al., 2021; Correa Magnus & Vasiliev, 2022; Koposov et al., 2023; Vasiliev, 2023; Fushimi et al., 2023; Watkins et al., 2024), $10-25\%$ the mass of the Milky Way itself (e.g., Bland-Hawthorn & Gerhard, 2016; Wang et al., 2020; Shen et al., 2022).
@@
Attempts to measure the LMC’s effects on the ‘bulk’ MW halo can be significantly confounded by the presence of undetected coherent substructure in the outer halo. The most prominent of these structures in the outer halo is the Sagittarius stream (Sgr), although 6D kinematics provide a powerful way to excise Sgr stars on the basis of their angular momenta (Johnson et al., 2020; Peñarrubia & Petersen, 2021). The LMC and SMC are themselves predicted to contribute stripped stars to the Galactic halo, forming a stellar counterpart to the well-known gaseous Magellanic Stream (MS; e.g., Besla et al., 2013). Zaritsky et al. (2020) discovered coherent debris in the H3 Survey (Conroy et al., 2019) at $\approx 50$ kpc, and argued that these might have a Magellanic origin. Chandra et al. (2023b) subsequently discovered a population of stars that appear to be stripped from the Clouds (the Magellanic Stellar Stream, or MSS), spanning the entire extent of the trailing MS from 50-100 kpc. The MSS and other debris could partially constitute the Pisces Plume overdensity, but there remains a wake-like structure in the outer halo once the MSS is removed (Chandra et al., 2023b; Amarante et al., 2024).
@@
Given the overlapping substructures present in the halo, full 6D kinematics are vital to disentangle coherent structures from the bulk MW halo. Recently, wide-field spectroscopic surveys — combined with astrometry from the Gaia space observatory — have enabled the kinematic detection of the LMC’s global influence on the MW halo. Erkal et al. (2021) compiled Galactocentric radial velocities of halo stars beyond 30 kpc, and found that stars in the northern Galactic hemisphere are systematically redshifted relative to those in the south, as expected for the LMC-induced reflex motion. The amplitude of their measured velocity dipole was broadly consistent with a $1.5\times 10^{11}\,M_{\odot}$ LMC. Stellar streams are also emerging as powerful tools to map the gravitational field of the Galaxy, including the LMC’s influence (e.g., Erkal et al., 2019b; Koposov et al., 2023; Lilleengen et al., 2023). Vasiliev et al. (2021) comprehensively modelled the formation of the Sgr stream in the presence of the LMC, finding that the ‘lurch’ of the inner MW towards the LMC is a crucial ingredient to reproduce the observed properties of Sgr.
@@
Petersen & Peñarrubia (2021, hereafter PP21) fit a physical model of the reflex velocity to the 6D phase-space positions of RGB and BHB stars beyond 40 kpc, and measured a global velocity vector $v_{\mathrm{travel}}=32\pm 4$ km s-1 with an apex (maximum velocity) direction that points towards a location along the past orbit of the LMC. While the amplitude of their measured signal is consistent with simulations of a $\gtrsim 10^{11}\,M_{\odot}$ LMC, the measured apex direction is somewhat different from most simulations of the MW-LMC interaction (Vasiliev, 2023). Whereas PP21 measure an apex direction in the opposite southern Galactic quadrant as the LMC today ($\ell_{\mathrm{apex}}>0^{\circ}$), most simulations predict that the velocity dipole should point towards a more recent point on the LMC’s orbit ($\ell_{\mathrm{apex}}<0^{\circ}$). More recent work by Yaaqib et al. (2024, hereafter YPP24) finds an increasing $v_{\mathrm{travel}}$ velocity towards larger distances using a sample spanning $6020<r_{\mathrm{gal}}/\mathrm{kpc}<$ 60, with an apex direction that varies by tens of degrees as a function of distance.
@@ 2.1 MagE Survey
Since 2022, we have been conducting a spectroscopic survey of the above selected giants with the Magellan Echellete Spectrograph (MagE; Marshall et al. 2008) on the 6.5m Magellan Baade Telescope at Las Campanas Observatory (PIs: Chandra & Naidu). The stars are typically between $17\lesssim G\lesssim 18.6$, and we utilize magnitude-dependent exposure times ranging from $10-40$ minutes per star to achieve signal-to-noise ratio (SNR) $\approx 10~{}\mathrm{pixel}^{-1}$ at $5100$. Spectra are reduced with a fully-automated pipeline [^1] https://github.com/vedantchandra/merlin/releases/tag/v0.2 built around the PypeIt utility (Prochaska et al., 2020).
@@
Figure 1: Validating the spectroscopic distances of stars in our MagE survey beyond 30 kpc. Black points show giant stars from the H3 Survey, which were fitted with a similar analysis pipeline as our MagE survey. The two colored points show mean distances of observed stars in the Crater 2 and Sextans dwarf galaxies, with their literature distances sourced from Torrealba et al. (2016) and Muñoz et al. (2018) respectively.
Figure 1 shows a cross-validation of the spectroscopic distance scale of our MagE survey. Apart from observing stars in common with the H3 Survey (Conroy et al., 2019), we observed giants from our parent giant catalog belonging to the Sextans ($86\pm 1$ kpc, Irwin et al. 1990; Muñoz et al. 2018) and Crater 2 ($117\pm 1$ kpc, Torrealba et al. 2016) dwarf galaxies. Figure 1 illustrates that we obtain reliable distance estimates out to $\gtrsim 100$ kpc. There is a $5\%$ systematic offset between the H3 and MagE spectroscopic distance scales, most likely caused by the MagE metallicity scale being $\approx 0.1$ dex lower than H3. Although we have not yet determined the cause of this offset, it is smaller than the typical statistical uncertainties on these distances. We have verified that artificially shrinking the MagE distances by $5\%$ to the H3 scale has a negligible effect on the results presented here.
As of May 2024, we have observed $400$ stars in this survey, of which $291$ are spectroscopically confirmed to lie beyond 50 kpc, and $106$ are beyond 100 kpc. This is already the largest dataset of stars with precise metallicities and 6D phase-space information beyond 100 kpc, and the only dataset of outer halo stars beyond $\approx 50$ kpc observed from the southern hemisphere. Our sample selection has a very low contamination rate from dwarfs ($\approx 10\%$), and the majority of the stars we observed have been confirmed to lie beyond $50$ kpc ($\approx 70\%$). All data from this survey will eventually be made publicly available, once our primary scientific projects have been completed.
@@ 2.2 H3 & SEGUE Data
The H3 Spectroscopic Survey (Conroy et al., 2019) has been conducting a spectroscopic survey of halo stars with the Hectochelle instrument (Szentgyorgyi et al., 2011) on the 6.5m MMT telescope at the Whipple Observatory in Arizona. We select stars observed by H3 up to January 2024 that have reliable stellar parameter fits from MINESweeper, and are not associated with cold substructures. Before applying any further selection cuts, there are $\approx 300$ giants beyond $d_{\mathrm{helio}}>50$ kpc from H3. Typical RV uncertainties for these data are of order $0.5-1$ km s-1 in absolute terms (Conroy et al., 2019; Cargile et al., 2020).
The Sloan Extension for Galactic Understanding and Exploration (SEGUE; Yanny et al. 2009) survey observed a quarter of a million stars with the low-resolution BOSS spectrograph as a part of the Sloan Digital Sky Survey (SDSS; York et al. 2000; Abazajian et al. 2009). We independently fit these spectra with the MINESweeper routine, from which $\approx 300$ stars with reliable parameters and distances $d_{\mathrm{helio}}>50$ kpc are used in this work. Due to the lower resolution of SEGUE data, these stars have typical RV uncertainties $\approx 5$ km s-1.
@@ 2.3 Sample Selection
Stars with unphysical radial or tangential velocities $>1000$ km s-1 are removed, as are stars within $1^{\circ}$ of known dwarf galaxies and globular clusters (McConnachie, 2012; Vasiliev & Baumgardt, 2021). The 13 stars identified by Chandra et al. (2023b) as belonging to the Magellanic Stellar Stream are excised. Finally, although stars in our MagE survey are selected to lie far from the LMC and SMC (see Figure 2), we search for any stars within $50^{\circ}$ of the LMC having proper motions and RVs consistent with belonging in an extended LMC halo. No stars confidently associated with the LMC are found.
Our final sample consists of $850$ field halo stars from SEGUE, H3, and MagE with $16040<r_{\mathrm{gal}}/\mathrm{kpc}<$ 160. The top panel of Figure 2 shows their on-sky spatial distribution in the outer halo, and the bottom panel shows their cumulative distance distributions. The bottom panel of Figure 2 illustrates the high efficiency of our MagE survey in pushing to the furthest Galactic outskirts, far outnumbering the number of $r_{\mathrm{gal}}\gtrsim 60$ kpc stars compared to the H3 and SEGUE surveys.
@@ 3.1 Radial Velocities
We begin by considering the overall radial velocity structure of the halo. Figure 3 shows the on-sky distribution of Galactocentric radial velocities $v_{\mathrm{GSR}}$ from our dataset, for stars with $r_{\mathrm{gal}}=50-100$ kpc. We use this limited distance range for this figure instead of our full $40-160$ kpc sample to minimize the effect of the different selection functions of the surveys combined here. As Figure 2 illustrates, most stars beyond $100$ kpc are from our MagE survey and lie only in the southern hemisphere. The $r_{\mathrm{gal}}=50-100$ kpc range ensures relatively even depth across the sky. In Appendix A, we further investigate the distance-dependence of these smoothed velocity maps, finding consistent behavior across all distances.
@@
It is immediately apparent from Figure 3 that the outer halo has a structured velocity distribution with variations across the sky. The Galactocentric radial velocity exhibits a net negative $\approx-12$ km s-1 signal. Furthermore, there is an unambiguous dipole signature in $v_{\mathrm{GSR}}$, with the southern halo systematically blueshifted, and the northern halo redshifted. Although this feature is most apparent in the smoothed map in the middle panel of Figure 3, it is visible even in the raw data in the top panel. This $v_{\mathrm{GSR}}$ dipole does not appear to be aligned purely in the north-south direction, but is rather tilted along an axis that approximately points towards the LMC. In Appendix A, we demonstrate that this observation is robust to the distance range considered, and rule out this effect being an artefact due to the more distant coverage of the MagE data.
@@
Figure 4 emphasizes what is apparent in the maps of Figure 3: these quadrants have drastically different radial velocities over a wide range of distances. Note that the bins uses to compute the average velocities in Figure 4 are entirely independent. In both quadrants, the absolute velocity grows as a function of distance. The 80-100 kpc bin in the LMC quadrant exhibits markedly negative radial velocities and warrants future investigation — although this could be due to unresolved substructure at these distances, it could also be a statistical fluctuation due to the small ($\approx 15$) number of stars in the bin. The observed behavior matches the MW+LMC simulations shown in Figure 4: the amplitude of the negative radial velocity signal in the southern quadrant reaches $\approx-45$ km s-1 at $\approx 130$ kpc, most consistent with the simulations containing a $>10^{\mathrm{11}}\,M_{\odot}$ LMC halo.

### v2

@@ 1 Introduction
The LMC and SMC (hereafter the ‘Clouds’) are by far the Milky Way’s (MW’s) most prominent satellite galaxies, and were likely the first external galaxies discovered (e.g., Dennefeld, 2020). There is growing evidence that the Clouds form a binary pair — with the LMC being 10$\approx$ 10$\times$ 10 × more massive than the SMC — and are on their first orbital infall towards the Milky Way (Kallivayalil et al., 2006; Besla et al., 2007, 2010; Kallivayalil et al., 2013, but see Vasiliev 2024 about the plausibility of a second infall scenario). Multiple independent measurements estimate the total mass of the LMC to be between $1.0-2.5\times 10^{11}\,M_{\odot}$ (Peñarrubia et al., 2016; Erkal et al., 2019b, 2020; Vasiliev et al., 2021; Shipp et al., 2021; Correa Magnus & Vasiliev, 2022; Koposov et al., 2023; Vasiliev, 2023; Fushimi et al., 2023; Watkins et al., 2024), $10-25\%$ the mass of the Milky Way itself (e.g., Bland-Hawthorn & Gerhard, 2016; Wang et al., 2020; Shen et al., 2022).
@@
Attempts to measure the LMC’s effects on the ‘bulk’ MW halo can be significantly confounded by the presence of undetected coherent substructure in the outer halo. The most prominent of these structures in the outer halo is the Sagittarius stream (Sgr), although 6D kinematics provide a powerful way to excise Sgr stars on the basis of their angular momenta (Johnson et al., 2020; Peñarrubia & Petersen, 2021). The LMC and SMC are themselves predicted to contribute stripped stars to the Galactic halo, forming a stellar counterpart to the well-known gaseous Magellanic Stream (MS; e.g., Besla et al., 2013; Petersen et al., 2022). Zaritsky et al. (2020) discovered coherent debris in the H3 Survey (Conroy et al., 2019) at $\approx 50$ kpc, and argued that these might have a Magellanic origin. Chandra et al. (2023b) subsequently discovered a population of stars that appear to be stripped from the Clouds (the Magellanic Stellar Stream, or MSS), spanning the entire extent of the trailing MS from 50-100 kpc. The MSS and other debris could partially constitute the Pisces Plume overdensity, but there remains a wake-like structure in the outer halo once the MSS is removed (Chandra et al., 2023b; Amarante et al., 2024).
@@
Given the overlapping substructures present in the halo, full 6D kinematics are vital to disentangle coherent structures from the bulk MW halo. Recently, wide-field spectroscopic surveys — combined with astrometry from the Gaia space observatory — have enabled the kinematic detection of the LMC’s global influence on the MW halo. Erkal et al. (2021) compiled Galactocentric radial velocities of halo stars beyond 30 kpc, and found that stars in the northern Galactic hemisphere are systematically redshifted relative to those in the south, as expected for the LMC-induced reflex motion. The amplitude of their measured velocity dipole was broadly consistent with a $1.5\times 10^{11}\,M_{\odot}$ LMC. Stellar streams are also emerging as powerful tools to map the gravitational field of the Galaxy, including the LMC’s influence (e.g., Erkal et al., 2019b; Koposov et al., 2023; Lilleengen et al., 2023; Brooks et al., 2024). Vasiliev et al. (2021) comprehensively modelled the formation of the Sgr stream in the presence of the LMC, finding that the motion of the inner MW towards the LMC is a crucial ingredient to reproduce the observed properties of Sgr.
@@
Petersen & Peñarrubia (2021, hereafter PP21) fit a physical model of the reflex velocity to the 6D phase-space positions of RGB and BHB stars beyond 40 kpc, and measured a global velocity vector $v_{\mathrm{travel}}=32\pm 4$ km s-1 with an apex (maximum velocity) direction that points towards a location along the past orbit of the LMC. While the amplitude of their measured signal is consistent with simulations of a $\gtrsim 10^{11}\,M_{\odot}$ LMC, the measured apex direction is somewhat different from most simulations of the MW-LMC interaction (Vasiliev, 2023). Whereas PP21 measure an apex direction in the opposite southern Galactic quadrant as the LMC today ($\ell_{\mathrm{apex}}>0^{\circ}$), many simulations predict that the velocity dipole should point towards a more recent point on the LMC’s orbit ($\ell_{\mathrm{apex}}<0^{\circ}$) (e.g., Garavito-Camargo et al., 2019; Vasiliev et al., 2021). More recent work by Yaaqib et al. (2024, hereafter YPP24) finds an increasing $v_{\mathrm{travel}}$ velocity towards larger distances using a sample spanning $6020<r_{\mathrm{gal}}/\mathrm{kpc}<$ 60, with an apex direction that varies by tens of degrees as a function of distance.
@@ 2.1 MagE Survey
We obtained spectroscopic observations of a subset of the above selected giants with the Magellan Echellete Spectrograph (MagE; Marshall et al. 2008) on the 6.5m Magellan Baade Telescope at Las Campanas Observatory (PIs: Chandra & Naidu). The stars are typically between $17\lesssim G\lesssim 18.6$, and we utilize magnitude-dependent exposure times ranging from $10-40$ minutes per star to achieve signal-to-noise ratio (SNR) $\approx 10~{}\mathrm{pixel}^{-1}$ at $5100$. Spectra are reduced with a fully-automated pipeline [^1] https://github.com/vedantchandra/merlin/releases/tag/v0.2 built around the PypeIt utility (Prochaska et al., 2020).
@@
Figure 1: Validating the spectroscopic distances of stars in our MagE survey beyond 30 kpc. Black points show giant stars from the H3 Survey, which were fitted with a similar analysis pipeline as our MagE survey. The two colored points show mean distances of observed stars in the Crater 2 and Sextans dwarf galaxies, with their literature distances sourced from Torrealba et al. (2016) and Muñoz et al. (2018) respectively. The lower panel shows the fractional residual, the difference between the MagE and literature distance divided by the literature distance.
Figure 1 shows a cross-validation of the spectroscopic distance scale of our MagE survey. Apart from observing stars in common with the H3 Survey (Conroy et al., 2019), we observed giants from our parent giant catalog belonging to the Sextans ($86\pm 1$ kpc, Irwin et al. 1990; Muñoz et al. 2018) and Crater 2 ($117\pm 1$ kpc, Torrealba et al. 2016) dwarf galaxies. Figure 1 illustrates that we obtain reliable distance estimates out to $\gtrsim 100$ kpc. The typical statistical distance uncertainties are smaller in the MagE dataset than H3 — for stars in common — since we aimed for higher SNR in the MagE survey, leading to more precise metallicities and consequently more precise isochrone distances. Using the scatter in the relation shown in Figure 1 as a guide, we adopt an uncertainty floor of 10% on our distance measurements. This is consistent with the standard deviation of the measured distances in Crater 2 and Sextans, which are both $\approx 10\%$.
There is a $5\%$ systematic offset between the H3 and MagE spectroscopic distance scales (see bottom panel of Figure 1), most likely caused by the MagE metallicity scale being $\approx 0.1$ dex lower than H3. Although we have not yet determined the cause of this offset, it is smaller than the typical statistical uncertainties on these distances. We have verified that artificially shrinking the MagE distances by $5\%$ to the H3 scale has a negligible effect on the results presented here. Furthermore, there is a $\approx 3$ km s-1 mean offset between MagE and H3 RVs of the same stars. We have likewise verified that shifting the MagE RVs to the H3 scale has a negligible effect on our results.
As of May 2024, we have observed $400$ stars in this survey, of which $291$ are spectroscopically confirmed to lie at heliocentric distances beyond 50 kpc, and $106$ are beyond 100 kpc. This is already the largest dataset of stars with precise metallicities and 6D phase-space information beyond 100 kpc, and by far the largest dataset of outer halo stars beyond $\approx 50$ kpc observed from the southern hemisphere. Our sample selection has a very low contamination rate from dwarfs ($\approx 10\%$). All data from this survey will eventually be made publicly available, once our primary scientific projects have been completed.
@@ 2.2 H3 & SEGUE Data
The H3 Spectroscopic Survey (Conroy et al., 2019) has been conducting a spectroscopic survey of halo stars with the Hectochelle instrument (Szentgyorgyi et al., 2011) on the 6.5m MMT telescope at the Whipple Observatory in Arizona. We select stars observed by H3 up to January 2024 that have reliable stellar parameter fits from MINESweeper, and are not associated with cold substructures (stars with on-sky positions and velocities consistent with membership in dwarf galaxies, star clusters, or known stellar streams). Before applying any further selection cuts, there are $\approx 300$ giants beyond $d_{\mathrm{helio}}>50$ kpc from H3. Typical RV uncertainties for these data are of order $0.5-1$ km s-1 in absolute terms (Conroy et al., 2019; Cargile et al., 2020).
The Sloan Extension for Galactic Understanding and Exploration (SEGUE; Yanny et al. 2009) survey observed a quarter of a million stars with the low-resolution BOSS spectrograph as a part of the Sloan Digital Sky Survey (SDSS; York et al. 2000; Abazajian et al. 2009). We independently fit these spectra with the MINESweeper routine, from which $\approx 300$ stars with reliable parameters and distances $d_{\mathrm{helio}}>50$ kpc are used in this work. Based on a comparison to literature parameters, we determine that SEGUE stars with SNR $<10$ per pixel do not produce reliable parameters, so we excise them from the dataset. Due to the low resolution of BOSS, the SEGUE stars have typical RV uncertainties $\approx 5$ km s-1.
@@ 2.3 Sample Selection
Stars with unphysical radial or tangential velocities $>1000$ km s-1 are removed, as are stars within $1^{\circ}$ of known dwarf galaxies and globular clusters (McConnachie, 2012; Vasiliev & Baumgardt, 2021). The 13 stars in our MagE dataset identified by Chandra et al. (2023b) as belonging to the Magellanic Stellar Stream are excised. Finally, although stars in our MagE survey are selected to lie far from the LMC and SMC (see Figure 2), we search for any stars within $50^{\circ}$ of the LMC having proper motions and RVs consistent with belonging in an extended LMC halo. No stars confidently associated with the LMC are found.
Our final sample consists of $851$ field halo stars from SEGUE, H3, and MagE with $16040<r_{\mathrm{gal}}/\mathrm{kpc}<$ 160. The top panel of Figure 2 shows their on-sky spatial distribution in the outer halo, and the bottom panel shows their cumulative distance distributions. The bottom panel of Figure 2 illustrates the high efficiency of our MagE survey in pushing to the furthest Galactic outskirts, far outnumbering the number of $r_{\mathrm{gal}}\gtrsim 60$ kpc stars compared to the H3 and SEGUE surveys. We adopt a systematic uncertainty floor of 1 km s-1 on RVs and 10% on distances. Since most stars have reported catalog uncertainties smaller than this, the median uncertainties in our final catalog are 1 km s-1 for RVs and 10% for distances. The median absolute magnitude of stars in our sample is $M_{\mathrm{G}}=-1.3$, the median metallicity is [Fe/H]$=-1.3$, and the median alpha abundance is [$\alpha$/Fe]$=+0.2$.
@@ 3.1 Radial Velocities
We begin by considering the overall radial velocity structure of the halo. Figure 3 shows the on-sky distribution of Galactocentric radial velocities $v_{\mathrm{GSR}}$ from our dataset, for stars with $r_{\mathrm{gal}}=50-100$ kpc. We use this limited distance range for this figure instead of our full $40-160$ kpc sample to minimize the effect of the different selection functions of the surveys combined here. As Figure 2 illustrates, most stars beyond $100$ kpc are from our MagE survey and lie only in the southern hemisphere. The $r_{\mathrm{gal}}=50-100$ kpc range ensures relatively even depth across the sky. In Appendix D, we further investigate the distance-dependence of these smoothed velocity maps, finding consistent behavior across all distances.
@@
It is immediately apparent from Figure 3 that the outer halo has a structured velocity distribution with variations across the sky. The Galactocentric radial velocity exhibits a net negative $\approx-12$ km s-1 signal. Furthermore, there is an unambiguous dipole signature in $v_{\mathrm{GSR}}$, with the southern halo systematically blueshifted, and the northern halo redshifted. Although this feature is most apparent in the smoothed map in the middle panel of Figure 3, it is visible even in the raw data in the top panel. This $v_{\mathrm{GSR}}$ dipole does not appear to be aligned purely in the north-south direction, but is rather tilted along an axis that approximately points towards the LMC. We emphasize that these are subtle mean trends in $v_{\mathrm{GSR}}$ brought out by averaging over large on-sky scales: the intrinsic radial velocity dispersion is 70--110 km s-1 depending on the sky location. In Appendix D, we demonstrate that this observation is robust to the distance range considered, and rule out this effect being an artifact due to the more distant coverage of the MagE data.
@@
Figure 4 emphasizes what is apparent in the maps of Figure 3: these quadrants have drastically different radial velocities over a wide range of distances. Note that the bins uses to compute the average velocities in Figure 4 are entirely independent. The 80-100 kpc bin in the LMC quadrant exhibits markedly negative radial velocities and warrants future investigation. This could be due to unresolved substructure at these distances, for example stars stripped from the LMC or SMC (e.g., Belokurov & Koposov, 2016; Chandra et al., 2023b). In Appendix D, we split the all-sky $v_{\mathrm{GSR}}$ maps by distance, showing that the overall behavior of these maps is consistent across distance bins. Furthermore, we have verified that removing the 26 stars in the discrepant bin of Figure 4 does not significantly alter the results from our 6D model fit in §$\S$ 4.2. The observed behavior matches the MW+LMC simulations shown in Figure 4: the amplitude of the negative radial velocity signal in the southern quadrant reaches $\approx-45$ km s-1 at $\approx 130$ kpc, most consistent with the simulations containing an LMC halo $\gtrsim 15\%$ the mass of the MW’s.